Operations Report – February 12th

Crew 205 Operations Report 12 – FEB – 2019

SOL: 2

Name of person filing report: Veronica Trivino

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on this evening by 1708 (12thFeb), Stopped by this morning 0730 (12thFeb)

Solar— SOC % – Turned off Gen (12th Feb) 33%, Turned on Gen (12th Feb) 69%

Diesel Reading – 100%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 77%, ~ 422 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 276.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01411709 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 49.6 hours

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Used

Hours: 73.7

Beginning charge: 100%

Ending charge: 40%

Currently charging: Yes

Notes on rovers: Spirit was accidentally left on after returned from EVA no.1. Engineer when checking readings in the evening turned it off. It was charged to 100% when checked by engineer.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? Not used

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Several suits with fan broken. Other suits had not finished charging by the time the second EVA preparation had begun, even though they were set to charge as soon as the crew from the first EVA came back. One of the suit chargers is not functional. More information was sent to Atila/Mission support in an email.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site